OpenStudy (anonymous):

Find the equation of the line through the points (-4, -2) and (-3, 5). y =______ x+ ___________

pooja195 (pooja195):

First find the slope \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

pooja195 (pooja195):

\[\LARGE \frac{5-(-2)}{-3-(-4)}=slope\] \[\LARGE \frac{5+2}{-3+4}=slope\]

OpenStudy (rhr12):

what about the value of b? @pooja195

pooja195 (pooja195):

You can do it once u find the slope

OpenStudy (anonymous):

it says "blank 1 and blank 2"

OpenStudy (anonymous):

what are the answers

pooja195 (pooja195):

Hi im sorry we dont GIVE answers we can GUIDE you.

OpenStudy (texaschic101):

in y = mx + b, the slope is in the m position and the b stands for the y intercept. first find the slope by using the formula above ^^. Then take 1 of your points...either one, and sub the x value in for x and the y value in for y and the slope (m) value in for m in y = mx + b...then solve for b. Do this and what do you get ?

pooja195 (pooja195):

Or u can use the point slope formula once u figure out the slope \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (texaschic101):

true ^^